# High Dimensional Distributed Gradient Descent with Arbitrary Number of Byzantine Attackers

### v2

@@ 3 Semi Verified Mean Estimation
Then we conduct spectral decomposition. If some eigenvalues of $V(S)$ are too large, then there should be some attacked samples which are moved in the directions of corresponding eigenvectors. This indicates that some samples need to be removed, based on their projection onto the subspace corresponds to these large eigenvalues. Based on these intuition, we conduct spectral decomposition of $V(S)$, such that $V(S)=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{T}$, with $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{d})$, $\lambda_{1}\geq\ldots\geq\lambda_{d}$. If there are $p$ eigenvalues larger than a certain threshold $\lambda_{c}$, i.e. $\lambda_{p}\geq\lambda_{c}$, then the algorithm removes some points to reduce the eigenvalue.
1
Input : Large corrupted dataset $S_{0}$, small clean dataset $A$
Output : Estimated mean $\hat{\mu}$
Parameter : $p$, $\lambda_{c}$
2 Initialize $S=S_{0}$;
3 $S=S\setminus\{i|\|\mathbf{Y}_{i}\|>N^{1/3}\}$;
4 while True do
5       Calculate sample mean $\mu(S)$ and sample covariance $V(S)$ using (4) and (5);
6       Conduct spectral decomposition of $V(S)$, such that $V(S)=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{T}$, with $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{d})$, $\lambda_{1}\geq\ldots\geq\lambda_{d}$;
7       if $\lambda_{p}\geq\lambda_{c}$ then
8             Let $\mathbf{P}=\mathbf{U}_{p}\mathbf{U}_{p}^{T}$, in which $\mathbf{U}_{p}\in\mathbb{R}^{d\times p}$ is the first $p$ columns of $\mathbf{U}$;
9             For each $i\in S$, calculate $\tau_{i}$ using (6);
10             $\tau_{\max}=\max_{i\in S}\tau_{i}$;
11             For each $i\in S$, remove $i$ from $S$ with probability $\tau_{i}/\tau_{\max}$;
12      else
13             break;
14       end if
15
16 end while
17Calculate $V(S)$, $\mathbf{P}$;
18 $\hat{\mu}=\mathbf{P}\mu(A)+(\mathbf{I}-\mathbf{P})\mu(S)$;
return $\hat{\mu}$;
Algorithm 1 Semi verified mean estimation
To determine the samples to remove, we assign to each sample $i\in S$ with a score
@@ 18 $\hat{\mu}=\mathbf{P}\mu(A)+(\mathbf{I}-\mathbf{P})\mu(S)$;
in which $\mathbf{P}=\mathbf{U}_{p}\mathbf{U}_{p}^{T}$, $\mathbf{U}_{p}\in\mathbb{R}^{d\times p}$ is the matrix of first $p$ columns of $\mathbf{U}$. $\mathbf{P}$ is a projection matrix to the space spanned by $p$ principal components. (6) can be understood as following. Firstly, we regularize the dataset by linear transformation with $V^{-\frac{1}{2}}(S)$. It is easy to show that samples after transformation have an identity covariance matrix:
$\frac{1}{|S|}\sum_{i\in S}\left[V^{-\frac{1}{2}}(S)(\mathbf{Y}_{i}-\mu(S))\right]\left[V^{-\frac{1}{2}}(S)(\mathbf{Y}_{i}-\mu(S))\right]^{T}=V^{-\frac{1}{2}}(S)V(S)V^{-\frac{1}{2}}(S)=\mathbf{I}_{d}.$ (7)
@@
After transformation, if a sample is far from the mean, i.e. $\|V^{-\frac{1}{2}}(S)(\mathbf{Y}_{i}-\mu(S))\|$ is large, then it is likely that this sample is attacked. However, as has been discussed earlier, it would be better to remove samples according to the projection into the subspace corresponds to large eigenvalues. Therefore, we conduct a projection with matrix $\mathbf{P}$, and then use the squared distance to the mean after the projection as the score. This is exactly $\tau_{i}$ in (6). Larger $\tau_{i}$ indicates higher confidence that sample $i$ is attacked. In the algorithm, sample $i$ is removed with probability $\tau_{i}/\tau_{\max}$, in which $\tau_{\max}=\max_{i\in S}\tau_{i}$ is the largest $\tau$ value among $S$. We would like to remark here that the random removal with probability $\tau_{i}/\tau_{\max}$ is only designed for the convenience of theoretical analysis. Practically, it may not be necessary to remove samples randomly. Instead, we can just remove $k$ samples with largest $\tau$ values.
@@
Sometimes $V(S)$ has some very small eigenvalues, thus accurate numerical computation of $V^{-\frac{1}{2}}(S)$ is hard. In this case, we can simplify the calculation of $\tau_{i}$ using (6). Consider that $\mathbf{P}=\mathbf{U}_{p}\mathbf{U}_{p}^{T}=\sum_{i=1}^{p}\mathbf{u}_{i}\mathbf{u}_{i}^{T}$, $V(S)=\sum_{i=1}^{d}\lambda_{i}\mathbf{u}_{i}\mathbf{u}_{i}^{T}$, in which $\mathbf{u}$ is the eigenvector corresponding to $\lambda_{i}$, hence $\mathbf{P}V^{-\frac{1}{2}}(S)=\sum_{i=1}^{p}\lambda_{i}^{-\frac{1}{2}}\mathbf{u}_{i}\mathbf{u}_{i}^{T}$. (6) can then be transformed to $\tau_{i}=\|\sum_{i=1}^{p}\lambda_{i}^{-\frac{1}{2}}\mathbf{u}_{i}\mathbf{u}_{i}^{T}(\mathbf{Y}_{i}-\mu(S))\|_{2}^{2}$. This format can avoid the numerical problem of calculating (6) directly.
The removal process is repeated, until $\lambda_{p}<\lambda_{c}$. Then we generate the final estimate $\hat{\mu}$ based on the intuition that the mean of $S$ is not accurate in directions with large eigenvalues, but accurate enough in other directions, since corrupted points that are far away from the mean in these directions have been removed. Therefore, for the directions with large eigenvalues, we use the reliable auxiliary data. For other directions, we just use $\mu(S)$. The resulting estimator is
$\hat{\mu}=\mathbf{P}\mu(A)+(\mathbf{I}-\mathbf{P})\mu(S).$ (8)
Furthermore, we would like to remark that the auxiliary clean data $A$ is only used in the last step (step 1). This is important since we need to use the assumption that the attacker has no knowledge of $A$, and ensure that the estimation error of $\mu(A)$ and $\mu(S)$ are independent.
4 Theoretical Analysis
@@
This section provides a theoretical analysis to the method proposed in Section 3. We use the following notations: $N_{A}$ is the size of auxiliary dataset, $N$ is the size of untrusted dataset $S_{0}$, i.e. $|S_{0}|=N$. $\alpha$ is the ratio of clean samples in $S_{0}$, i.e. $|S_{0}^{*}|\geq\alpha N$. For two matrices $\mathbf{A}$ and $\mathbf{B}$, $\mathbf{A}\preceq\mathbf{B}$ if $\mathbf{B}-\mathbf{A}$ is positive semidefinite. $a\lesssim b$ if $a\leq Cb$ for some constant $C$ that does not depend on $N_{A}$, $N$, $d$ and $\alpha$.
@@
We begin with the following assumption, which requires the boundedness of covariance matrix:
@@
Denote $\mu^{*}=\mathbb{E}[\mathbf{X}]$ and $V^{*}=\mathbb{E}[(\mathbf{X}-\mu^{*})(\mathbf{X}-\mu^{*})^{T}]$ as the mean and covariance matrix, respectively, in which $\mathbf{X}\sim D$. Then
$V^{*}\preceq\sigma^{2}\mathbf{I}_{d}.$ (9)
Right now we only assume that the distribution $D$ has bounded covariance matrix. Under stronger tail assumptions, such as boundedness of higher order moments, it is possible to further improve the convergence [42, 43].
4.1 Upper Bound
@@
Under additive contamination model, if Assumption 1 and 2 hold, and parameters $p$, $\lambda_{c}$ satisfy
@@
$\lambda_{c}$ $>$ $32\sigma^{2}\left(1+\frac{2d}{\alpha N}\right),$ (11)
@@
$\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\leq\frac{3\sigma^{2}p}{N_{A}}+\frac{15\lambda_{c}}{2\alpha}+\delta_{N},$ (12)
in which $\delta_{N}$ decays faster than any polynomial of $N$.
Proof.
The proof is shown in Section A in the appendix. ∎
Theorem 1 shows that, if $N\gtrsim d/\alpha$, $\lambda_{c}\sim\sigma^{2}$, $1/\alpha\lesssim p\lesssim N_{A}/\alpha$, then $\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\lesssim\sigma^{2}/\alpha$, and $\mathbb{E}[\|\hat{\mu}-\mu^{*}\|]\lesssim\sigma\alpha^{-1/2}$.
Algorithm 1 is a direct method for semi-verified mean estimation. This task can also be solved indirectly by converting it to list-decodable mean estimation problem [28]. Now we compare our direct method with the indirect approach:
  • Our direct method has better sample efficiency. In particular, it requires less auxiliary clean samples. Even if $N_{A}=1$, we can let $p\sim 1/\alpha$, then $\mathbb{E}[\|\hat{\mu}-\mu^{*}\|]\lesssim\sigma\alpha^{-1/2}$. However, for indirect method, a single auxiliary sample is not enough. According to discussions in [28] and [38], the list-decodable mean estimation methods generate a hypotheses list $\mathcal{L}$ with length $O(1/\alpha)$ whose minimum distance to the ground truth is $O(\sigma\alpha^{-1/2})$. After that, at least $O(\ln(1/\alpha))$ additional clean samples are needed to pick the optimal one from $\mathcal{L}$.
  • If $N_{A}>1$, then our method does not need precise knowledge of $\alpha$. Previous list decodable mean estimation methods [28, 37, 29, 42, 38] requires the knowledge of $\alpha$, which is usually not practical. These methods can use a conservative estimate of $\alpha$ that is lower than its truth, but the accuracy will be sacrificed. However, our method only requires $1/\alpha\lesssim p\lesssim N_{A}/\alpha$, which is a broad range if $N_{A}>1$, while the other parameter $\lambda_{c}$ does not depend on $\alpha$. As a result, our method is easier to use since the precise knowledge of $\alpha$ is no longer necessary.
@@
Under strong contamination model, if Assumption 1 and 2 hold, $p$ satisfies (10), and $\lambda_{c}$ satisfies
$\lambda_{c}>\frac{8\sigma^{2}}{\alpha}\left(1+\sqrt{\frac{16d\ln^{2}N}{3\alpha N}}\right)^{2},$ (13)
@@
$\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\leq\frac{3\sigma^{2}p}{N_{A}}+\frac{15\lambda_{c}}{2\alpha}+\delta_{N},$ (14)
in which $\delta_{N}$ decays faster than any polynomial of $N$.
Theorem 2 is about the error rate under strong contamination model. Although (14) appears to be the same as (12) for additive contamination model, the minimum value of $\lambda_{c}$ is different between these two models. If $N/\ln^{2}N\gtrsim d$, ignoring constant factor, (13) is larger than (11) with a roughly $1/\alpha$ factor. Hence, under strong contamination model, the performance is worse than in additive model. In particular, given $N/\ln^{2}N\gtrsim d/\alpha$, $\lambda_{c}\sim\sigma^{2}/\alpha$, $1/\alpha\lesssim p\lesssim N_{A}/\alpha$, then $\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\lesssim\sigma^{2}/\alpha^{2}$, and $\mathbb{E}[\|\hat{\mu}-\mu^{*}\|]\lesssim\sigma/\alpha$.
Compared with the additive model, under strong contamination model, the estimation error is larger up to a $\alpha^{-\frac{1}{2}}$ factor. As has been discussed earlier, under additive model, good samples are mixed with adversarial samples. The distribution of good samples is not modified by the attacker. On the contrary, under strong contamination model, the attacker can inspect all samples in $S_{0}$ and maliciously select $(1-\alpha)N$ of them to modify their values. As a result, the distribution of remaining good samples can become far away from $D$.
4.2 Minimax lower bound
@@
in which $\mathcal{F}$ is the set of all distributions satisfying Assumption 2. $\pi_{A}(\alpha)$ is the policy of attacker, which maps $\mathbf{X}_{i}$ for $i\in S_{0}$ to $\mathbf{Y}_{i}$, $i\in S_{0}$. In particular, it picks $\lceil\alpha N\rceil$ samples randomly, and let $\mathbf{Y}_{i}=\mathbf{X}_{i}$ for these samples. For other samples, $\mathbf{Y}_{i}$ are arbitrary. $\hat{\mu}$ is the estimator, which can be any function of $\mathbf{Y}_{i}$ for $i\in S_{0}$ and $\mathbf{X}_{i}$ for $i\in A$. Similarly, under strong contamination model, the minimax risk is defined as
@@
If
$N_{A}\leq\frac{\ln\frac{d}{4}}{2\alpha\left(\ln\frac{2}{\alpha}+1\right)},$ (17)
then with probability at least $1/2-\exp[-(\ln 2-1/2)N\alpha]$,
$R_{A}(\alpha)$$\geq$ $\frac{\sigma}{\sqrt{2\alpha}},$ (18)
$R_{S}(\alpha)$$\geq$ $\frac{\sigma}{\sqrt{2}\alpha}.$ (19)
The proof of Theorem 3 is provided in section C. The upper bound matches the minimax lower bound. Such results indicate that the error rates of Algorithm 1 are optimal. In other words, it is impossible to further improve the bounds in Theorem 1 and 2 in general.
5 Application in Distributed Learning under Byzantine Attack
In this section, we formalize the problem of distributed learning under Byzantine failures. Our formalization follows [13, 14, 19].
Suppose that all training samples are i.i.d following a common distribution $Q$. Let $f(\mathbf{w},\mathbf{z})$ be a loss function of a parameter vector $\mathbf{w}\in\Omega\subseteq\mathbb{R}^{d}$, in which $\Omega$ is the parameter space. Correspondingly, define the population risk function
$F(\mathbf{w})=\mathbb{E}[f(\mathbf{w},\mathbf{Z})],$ (20)
in which $\mathbf{Z}$ follows distribution $Q$. Our goal is to learn the model with the minimizer of population risk function $F$:
$\mathbf{w}^{*}=\underset{\mathbf{w}\in\Omega}{\arg\min}F(\mathbf{w}).$ (21)
We consider the following model. Suppose there is a master machine $W_{0}$ and $m$ working machines, $W_{1},\ldots,W_{m}$. A clean dataset $A=\{\mathbf{Z}_{01},\ldots,\mathbf{Z}_{0N_{A}}\}$ is stored in the master, with size $|A|=N_{A}$. Each worker machine $W_{i}$ has $n$ samples $\mathbf{Z}_{i1},\ldots,\mathbf{Z}_{in}$. All samples stored in both master and working machines follow distribution $Q$. Under this setting, $\mathbf{w}^{*}$ can be learned via gradient descent, in which gradient can be estimated using sample gradients of each working machine. However, a large fraction of working machines are Byzantine. Denote $\mathcal{B}$ as the set of Byzantine machines. It is guaranteed that at least $\alpha m$ working machines are not attacked, i.e. $|\mathcal{B}|\leq(1-\alpha)m$.
In each iteration, the master machine broadcasts current parameter $\mathbf{w}_{t}$ to all working machines. Then these working machines calculate the gradient using the local dataset:
$\mathbf{X}_{i}=\frac{1}{n}\sum_{j=1}^{n}\nabla f(\mathbf{w},\mathbf{Z}_{ij}).$ (22)
For any normal working machines, the master can receive $\mathbf{X}_{i}$ exactly. On the contrary, Byzantine machines can send arbitrary messages determined by the attacker. Denote $\mathbf{Y}_{i}\in\mathbb{R}^{d}$ as the vector received by the master, then
$\mathbf{Y}_{i}=\left\{\begin{array}[]{ccc}\mathbf{X}_{i}&\text{if}&i\notin\mathcal{B}\\ \star&\text{if}&i\in\mathcal{B}.\end{array}\right.$ (25)
We use the semi-verified mean estimation algorithm (Algorithm 1) to estimate the gradient vector. The parameter $\mathbf{w}$ is then updated using this estimated gradient. The detailed steps are shown in Algorithm 2.
1
Input : Master machine $W_{0}$, working machines $W_{1},\ldots,W_{m}$
Output : Estimated weight $\hat{\mathbf{w}}$
Parameter : Initial weight parameter $\mathbf{w}_{0}\in\Omega$, step length $\eta$, algorithm parameters $p$, $\lambda_{c}$
2 for $t=0,1,\ldots,T-1$ do
3       Master machine: broadcast current parameter $\mathbf{w}_{t}$ to all working machines;
4       for $i\in[m]$ in parallel do
5             Worker machine $i$: compute local gradient $\mathbf{X}_{i}$ using (22);
6             if $i$ is normal machine then
7                   send $\mathbf{X}_{i}$ to master;
8            else
9                   send arbitrary $d$ dimensional vector to master;
10             end if
11
12       end for
13      Master machine: Receive $\mathbf{Y}_{i}$, $i=1,\ldots,m$ from each working machine;
14       Calculate aggregated gradient $g(\mathbf{w}_{t})$ using Algorithm 1, in which $S_{0}=\{\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m}\}$, $A$ is the clean dataset stored in the master, with parameter $p,\lambda_{c}$;
15       Update parameter $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\eta g(\mathbf{w}_{t})$;
16 end for
Algorithm 2 Robust Distributed Gradient Descent
We then provide a theoretical analysis to the distributed learning method in Algorithm 2, which uses our new semi-verified mean estimation method as the aggregator function. Our analysis is based on the following assumption.
@@ 15       Update parameter $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\eta g(\mathbf{w}_{t})$;
$\underset{\mathbf{v}:\|\mathbf{v}\|_{2}=1}{\sup}\mathbb{E}\left[\exp\left(\lambda\mathbf{v}^{T}\left(\nabla f(\mathbf{w},\mathbf{Z})-\nabla F(\mathbf{w})\right)\right)\right]\leq e^{\frac{1}{2}\sigma^{2}\lambda^{2}},\forall|\lambda|\leq\frac{1}{\sigma};$ (26)
(b) $F(\mathbf{w})$ is $\alpha$-strong convex and $L$-smooth in $\mathbf{w}$.
(a) is more restrictive than Assumption 2. (a) requires the distribution of gradient values to be sub-exponential, while the latter only requires bounded eigenvalues of covariance matrix. Such strengthened assumption is made only for theoretical completeness. For (b), despite that our theoretical results are derived under the assumption that $F$ is strong convex, similar to [14], our analysis can be easily generalized to the case with non-strong convex and nonconvex functions.
@@
Suppose that the following conditions are satisfied:
(1) Assumption 1 and 3 hold;
(2) $p>8/\alpha$;
(3) $\lambda_{c}$ satisfies
@@
$\lambda_{c}>\frac{8\sigma^{2}}{n\alpha}\left(1+\sqrt{\frac{16d\ln^{2}m}{3\alpha m}}\right)^{2}.$ (28)
under strong contamination model;
(4) $\eta\leq 1/L$;
(5) The size of auxiliary clean dataset satisfies
$N_{A}\geq 2\ln\frac{mT}{\delta}.$ (29)
Then with probability at least $1-\delta-Te^{-\frac{1}{64}\alpha N}-4Tme^{-\frac{1}{16}\lambda_{c}m\alpha^{2}m^{\frac{1}{3}}\epsilon^{2}}$,
$\|\hat{\mathbf{w}}-\mathbf{w}^{*}\|\leq(1-\rho)^{T}\|\mathbf{w}_{0}-\mathbf{w}^{*}\|_{2}+\frac{\eta\Delta}{\rho},$ (30)
in which $\hat{\mathbf{w}}=\mathbf{w}_{T}$ is the updated weight after $T$ iterations, and
$\rho=\frac{1}{2}\eta\alpha,$ (31)

### v3

@@ Introduction
If some eigenvalues of $V(S)$ are too large, then there should be some attacked samples which are moved in the directions of corresponding eigenvectors by the adversary. Motivated by this intuition, we conduct spectral decomposition of $V(S)$, such that $V(S)=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{T}$, with $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{d})$, $\lambda_{1}\geq\ldots\geq\lambda_{d}$. If there are $p$ eigenvalues larger than a certain threshold $\lambda_{c}$, i.e. $\lambda_{p}\geq\lambda_{c}$, then the algorithm removes some points to reduce the eigenvalue. In particular, each sample $i$ is assigned a score
@@
in which $\mathbf{P}=\mathbf{U}_{p}\mathbf{U}_{p}^{T}$, with $\mathbf{U}_{p}\in\mathbb{R}^{d\times p}$ being the matrix of first $p$ columns of $\mathbf{U}$. $\mathbf{P}$ is a projection matrix to the space spanned by $p$ principal components.
We now give an intuitive explanation of (10). We first regularize the dataset by linear transformation with $V^{-\frac{1}{2}}(S)$. It is easy to show that samples after transformation have an identity covariance matrix:
$\hskip-28.45274pt\frac{1}{|S|}\sum_{i\in S}\left[V^{-\frac{1}{2}}(S)(\mathbf{Y}_{i}-\mu(S))\right]\left[V^{-\frac{1}{2}}(S)(\mathbf{Y}_{i}-\mu(S))\right]^{T}$ (11)
$=$ $V^{-\frac{1}{2}}(S)V(S)V^{-\frac{1}{2}}(S)=\mathbf{I}_{d}.$
@@
After transformation, if a sample is far from the mean, i.e. $\|V^{-\frac{1}{2}}(S)(\mathbf{Y}_{i}-\mu(S))\|$ is large, then it is likely that this sample is attacked. As has been discussed earlier, it would be better to filter samples in directions corresponding to large eigenvalues. Therefore, (10) incorporates the projection matrix $\mathbf{P}$. Larger $\tau_{i}$ indicates higher confidence that sample $i$ is attacked. With this intuition, sample $i$ is removed with probability $\tau_{i}/\tau_{\max}$, where $\tau_{\max}=\max_{i\in S}\tau_{i}$. This randomized removal is mainly for theoretical convenience; in practice, one may simply remove the $k$ samples with the largest $\tau$ values. The removal process continues until $\lambda_{p}<\lambda_{c}$.
@@
Sometimes $V(S)$ has some very small eigenvalues, thus accurate numerical computation of $V^{-\frac{1}{2}}(S)$ is hard. In this case, the calculation can be simplified to $\tau_{i}=\|\sum_{j=1}^{p}\lambda_{j}^{-\frac{1}{2}}\mathbf{u}_{j}\mathbf{u}_{j}^{T}(\mathbf{Y}_{i}-\mu(S))\|_{2}^{2}$. This format can avoid the numerical problem of calculating (10) directly.
(3) Calculate $V(S)$ and $\mathbf{P}$ again. At step 15, after iterative removal, the algorithm calculates the projection matrix using step 7 in Algorithm 2. This is illustrated in Figure 1(c), in which the blue arrow represents the projection operation.
(4) Calculate $\hat{\mu}$. The semi-verified mean estimator is
$\hat{\mu}=\mathbf{P}\mathbf{X}_{0}+(\mathbf{I}-\mathbf{P})\mu(S).$ (12)
As shown in Figure 1(d), both $\mathbf{X}_{0}$ and $V(S)$ are relatively far from the ground truth $\mu^{*}$. However, with appropriate projection, the final estimate with (12) is much closer to $\mu^{*}$.
Theoretical Analysis
@@
This section provides theoretical analysis about the semi-verified mean estimation method in Algorithm 2. From now on, we use the following notations: For two matrices $\mathbf{A}$ and $\mathbf{B}$, $\mathbf{A}\preceq\mathbf{B}$ if $\mathbf{B}-\mathbf{A}$ is positive semidefinite. $a\lesssim b$ if $a\leq Cb$ for some constant $C$ that does not depend on $N_{A}$, $m$, $n$, $d$ and $\alpha$. $\operatorname{Cov}[\mathbf{U}]$ denotes the covariance matrix of random vector $\mathbf{U}$. The proofs are shown in the appendix.
@@
Our analysis begins with the following assumption, which requires the boundedness of covariance matrix:
@@
$V_{0}\preceq\sigma^{2}\mathbf{I}_{d}$, in which $V_{0}$ is the covariance matrix of $\nabla f(\mathbf{w}_{t},\mathbf{Z}_{ij})$.
According to (7), we have $V^{*}\preceq(\sigma^{2}/n)\mathbf{I}_{d}$ and $V_{A}^{*}\preceq(\sigma^{2}/N_{A})\mathbf{I}_{d}$.
Upper Bound
@@
Under additive contamination model, if Assumption 1 hold, and parameters $p$, $\lambda_{c}$ in Algorithm 2 satisfy
@@
$\lambda_{c}$ $>$ $32\frac{\sigma^{2}}{n}\left(1+\frac{2d}{\alpha m}\right),$ (14)
@@
$\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\leq\frac{3\sigma^{2}p}{N_{A}}+\frac{15\lambda_{c}}{2\alpha}+\delta_{m},$ (15)
in which $\delta_{m}$ decays faster than any polynomial of $m$.
With $m\gtrsim d/\alpha$, from (14), we can let $\lambda_{c}\sim\sigma^{2}/n$, then (15) becomes
$\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\lesssim\sigma^{2}\left(\frac{p}{N_{A}}+\frac{1}{\alpha n}\right).$ (16)
We now discuss the selection of parameter $p$. Since (13) requires $p\gtrsim 1/\alpha$, we discuss two cases. If $N_{A}\lesssim n$, according to (16), let $p\sim 1/\alpha$, then $\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\lesssim\sigma^{2}/(\alpha N_{A})$. If $N_{A}\gtrsim n$, according to (16), as long as $p$ satisfies $1/\alpha\lesssim p\lesssim N_{A}/(\alpha n)$, then $\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\lesssim\sigma^{2}/(\alpha n)$. With this selection rule of $p$, (16) becomes
$\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}\right]\lesssim\sigma\alpha^{-\frac{1}{2}}\left(\frac{1}{\sqrt{N_{A}}}+\frac{1}{\sqrt{n}}\right).$ (17)
@@
Under strong contamination model, if Assumption 1 hold, $p$ satisfies (13), and $\lambda_{c}$ satisfies
$\lambda_{c}>\frac{8\sigma^{2}}{\alpha n}\left(1+\sqrt{\frac{16d\ln^{2}m}{3\alpha m}}\right)^{2},$ (18)
@@
$\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\leq\frac{3\sigma^{2}p}{N_{A}}+\frac{15\lambda_{c}}{2\alpha}+\delta_{m},$ (19)
in which $\delta_{m}$ decays faster than any polynomial of $m$.
Despite that (19) appears to be the same as (15) for additive contamination model, the minimum value of $\lambda_{c}$ is different between these two models. With $m/\ln^{2}m\gtrsim d/\alpha$, from (18), we let $\lambda_{c}\sim\sigma^{2}/(\alpha n)$, then (19) becomes
$\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}^{2}\right]\lesssim\sigma^{2}\left(\frac{p}{N_{A}}+\frac{1}{\alpha^{2}n}\right).$ (20)
If $N_{A}\lesssim\alpha n$, according to (20), let $p\sim 1/\alpha$. Otherwise, just need to ensure that $1/\alpha\lesssim p\lesssim N_{A}/(\alpha^{2}n)$. Then (20) becomes
$\mathbb{E}\left[\|\hat{\mu}-\mu^{*}\|_{2}\right]\lesssim\sigma\alpha^{-\frac{1}{2}}\left(\frac{1}{\sqrt{N_{A}}}+\frac{1}{\sqrt{\alpha n}}\right).$ (21)
Traditional methods, such as Zeno (xie2019zeno; xie2020zeno++) has an error of $\sigma\alpha^{-1/2}(\sqrt{d/N_{A}}+\sqrt{d/n})$. In contrast, our approach—as shown in (17) and (21)—does not suffer from the $\sqrt{d}$ dependence, making it significantly more effective for high-dimensional mean estimation. Depending on whether we are assuming additive or strong contamination, the results are slightly different ($1/\sqrt{n}$ vs $1/\sqrt{\alpha n}$). As discussed earlier, strong contamination model gives the attacker more room for manipulation, since it allows the attacker to pick samples arbitrarily instead of randomly, thus the estimation error is larger.
Remark 1.
(diakonikolas2021list) has proposed an efficient method for list-decodable mean estimation, which generates a list of $O(1/\alpha)$ hypotheses whose minimum distance to $\mu^{*}$ is $O(1/\sqrt{\alpha})$. As is discussed in (charikar2017learning), list-decodable method can be used in semi-verified mean estimation. Compared with this indirect approach, our new method requires less number of auxiliary clean samples, and the parameter selection is more flexible. Detailed comparisons can be found in the appendix.
Minimax lower bound
@@
in which $\mathcal{F}$ is the set of all distributions satisfying Assumption 1, as well as (7), which implies that $\operatorname{Cov}[\mathbf{X}_{i}]\preceq(\sigma^{2}/n)\mathbf{I}$ and $\operatorname{Cov}[\mathbf{X}_{0}]\preceq(\sigma^{2}/N_{A})\mathbf{I}$. $\pi_{A}(\alpha)$ is the policy of attacker, which maps $\mathbf{X}_{i}$ for $i\in S_{0}$ to $\mathbf{Y}_{i}$, $i\in S_{0}$. In particular, it picks $\lceil\alpha N\rceil$ samples randomly, and let $\mathbf{Y}_{i}=\mathbf{X}_{i}$ for these samples. For other samples, $\mathbf{Y}_{i}$ are arbitrary. $\hat{\mu}$ is the estimator, which is a function of $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m},\mathbf{X}_{0}$.
Similarly, under strong contamination model, the minimax risk is defined as
@@
If $\left\lceil N_{A}/n\right\rceil\leq\ln\frac{d}{4}/\left(4\beta\left(\ln\frac{1}{\beta}+1\right)\right)$, then with probability at least $1/2-\exp[-(\ln 2-1/2)m\alpha]$,
$R_{A}(\alpha)$ $\geq$ $\frac{\sigma}{2\sqrt{2\alpha}}\left(\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{N_{A}}}\right),$ (24)
$R_{S}(\alpha)$ $\geq$ $\frac{\sigma}{2\sqrt{2\alpha}}\left(\frac{1}{\sqrt{\alpha n}}+\frac{1}{\sqrt{N_{A}}}\right).$ (25)
The upper bound matches the minimax lower bound up to constant factors. Such results indicate that the error rates of Algorithm 2 are optimal. In other words, it is impossible to further improve the bounds in Theorem 1 and 2 in general.
Application in Distributed Learning under Byzantine Attack
Based on the analysis of semi-verified mean estimation in the previous section, we now analyze the distributed learning method in Algorithm 1, which uses our new semi-verified mean estimation method as the aggregator function. Our analysis is based on the following assumption.
@@
$\underset{\mathbf{v}:\|\mathbf{v}\|_{2}=1}{\sup}\mathbb{E}\left[e^{\lambda\mathbf{v}^{T}\left(\nabla f(\mathbf{w},\mathbf{Z})-\nabla F(\mathbf{w})\right)}\right]\leq e^{\frac{1}{2}\sigma^{2}\lambda^{2}},$ (26)
for all $\lambda$ with $|\lambda|\leq 1/\sigma$;
(b) $F(\mathbf{w})$ is $\mu$-strong convex and $L$-smooth in $\mathbf{w}$.
(a) is more restrictive than Assumption 1. (a) requires the distribution of gradient values to be sub-exponential, while the latter only requires bounded eigenvalues of covariance matrix. Such strengthened assumption is made only for theoretical completeness. For (b), despite that our theoretical results are derived under the assumption that $F$ is strong convex, similar to (yin2018byzantine), our analysis can be easily generalized to the case with non-strong convex and nonconvex functions.
@@
Suppose that the following conditions are satisfied: (1) Assumption 2 hold; (2) $p>8/\alpha$; (3) $\lambda_{c}$ satisfies
@@
$\lambda_{c}>\frac{8\sigma^{2}}{\alpha n}\left(1+\sqrt{\frac{16d\ln^{2}m}{3\alpha m}}\right)^{2}$ (28)
under strong contamination model; (4) $\eta\leq 1/L$; (5) The size of auxiliary clean dataset satisfies $N_{A}\geq 2\ln(mT/\delta)$.
Then with probability at least $1-\delta-Te^{-\frac{1}{64}\alpha m}-4Tme^{-\frac{1}{16}\lambda_{c}m\alpha^{2}m^{\frac{1}{3}}\epsilon^{2}}$,
$\|\hat{\mathbf{w}}-\mathbf{w}^{*}\|\leq(1-\rho)^{T}\|\mathbf{w}_{0}-\mathbf{w}^{*}\|_{2}+\frac{2\Delta}{\mu},$ (29)
in which $\hat{\mathbf{w}}=\mathbf{w}_{T}$ is the updated weight after $T$ iterations, and $\rho=\eta\mu/2$,